OpenStudy (anonymous):

Find a fourth degree polynomial function having only real coefficients with 3, 1, and (-1+3i) as zeros and f(2)=-36?

OpenStudy (aum):

Start with the complex factor: x = -1+3i (x+1) = 3i Square: (x+1)^2 = (3i)^2 x^2+2x+1 = -9 x^2+2x+10 = 0 The other two factors are (x-3) and (x-1) multiply them all and expand: (x^2+2x+10) * (x-3) * (x-1)

OpenStudy (aum):

The fourth degree polynomial is: f(x) = a * (x^2+2x+10) * (x-3) * (x-1) Put x = 2 to evaluate f(2). Equate it to -36 and solve for a.

OpenStudy (anonymous):

I did just as you had stated, and ended up with x^4-2x^3+5x^2-34x+30 When I plugged in f(2) I ended up with -18. Can you tell me where I went wrong?

OpenStudy (aum):

f(x) = a * (x^2+2x+10) * (x-3) * (x-1) where 'a' is a constant to be determined. Find f(2): f(2) = a* (4+4+10)(-1)(1) = a * (-18) But we are given f(2) = -36 Therefore, -18a = -36 a = 2. So f(x) = 2(x^4-2x^3+5x^2-34x+30) or f(x) = 2x^4 - 4x^3 + 10x^2 - 68x + 60